OpenStudy (anonymous):

The acceleration of an object can be described by the equation a=2d/t^2, where a is acceleration d is the distance, and t is time. if an object accelerates at a rate of 2(m/s^2) for 10 meters what is the total time elapsed?

OpenStudy (jim766):

\[A = \frac{ 2d }{ t^2 }\] fill in what the gave you \[2=\frac{ 2(10) }{ t^2 }\] need to get t alone, mult both sides by t^2

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

what both sides?

OpenStudy (jim766):

right what is 2 times t^2?

OpenStudy (anonymous):

im confused will you explain it a little more.

OpenStudy (anonymous):

i dont know how to times that.

OpenStudy (jim766):

\[2(t^2)=\frac{ 20(t^2) }{ t^2 }\] the 2 t^2 on the right cancel out \[2t^2= 20\]

OpenStudy (jim766):

now divide both sides by 2

OpenStudy (anonymous):

10?

OpenStudy (jim766):

t^2 = 10

OpenStudy (jim766):

to change the t^2 to just t, take the square root of both sides

OpenStudy (jim766):

\[t=\sqrt{10}\]

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

thank you :D

OpenStudy (jim766):

is that good?

OpenStudy (anonymous):

yes can u help with another simliar one?

OpenStudy (jim766):

I'll try

OpenStudy (anonymous):

The acceleration of an object can be described by the equation a=2d/t^2, where a is acceleration d is the distance, and t is time. suppose an object accelerates for a distance of 15 meters. which of the following equations relates the time to the acceleration?

OpenStudy (jim766):

this time the only thing that give you is the distance. just sub it in

OpenStudy (anonymous):

hmm what do you mean?

OpenStudy (jim766):

\[a =\frac{ 2(15) }{ t^2 }\] they tell you 15 is the distance...replace the d in the formula with 15

OpenStudy (anonymous):

okay, now what?

OpenStudy (jim766):

simplify 2*15

OpenStudy (anonymous):

30

OpenStudy (jim766):

\[a=\frac{ 30 }{ t^2 }\]

OpenStudy (anonymous):

yes. thank you :))

OpenStudy (jim766):

yw